### v1

# BugCraft: End-to-End Crash Bug Reproduction Using LLM Agents in Minecraft

Abstract.
Context: Reproducing game bugs, in our case crash bugs in continuously evolving games like Minecraft, is a notoriously manual, time-consuming, and challenging process to automate. Despite the success of LLM-driven bug reproduction in other software domains, games, with their complex interactive environments, remain largely unaddressed.
Objective: This paper introduces BugCraft, a novel end-to-end framework designed to automate the reproduction of crash bugs in Minecraft directly from user-submitted bug reports, addressing the critical gap in automated game bug reproduction.
Method: BugCraft employs a two-stage approach: first, a Step Synthesizer leverages LLMs and Minecraft Wiki knowledge to transform bug reports into high-quality, structured steps to reproduce (S2R). Second, an Action Model, powered by a vision-based LLM agent (GPT-4o) and a custom macro API, executes these S2R steps within Minecraft to trigger the reported crash. To facilitate evaluation, we introduce BugCraft-Bench, a curated dataset of Minecraft crash bug reports.
Results: Evaluated on BugCraft-Bench, our framework successfully reproduced 30.23% of crash bugs end-to-end. The Step Synthesizer demonstrated a 66.28% accuracy in generating correct bug reproduction plans, highlighting its effectiveness in interpreting and structuring bug report information.
Conclusion: BugCraft demonstrates the feasibility of automated reproduction of crash bugs in complex game environments using LLMs, opening promising avenues for game testing and development. The framework and the BugCraft-Bench dataset pave the way for future research in automated game bug analysis and hold potential for generalization to other interactive game platforms. Finally, we make our code open at https://bugcraft2025.github.io/
Automated Bug Reproduction, Vision-Based Agent, Large Language Model, Crash Bugs, Minecraft, Bug Tracking
copyright: ccjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/24/06ccs: Software and its engineering Software testing and debuggingccs: Software and its engineering Maintaining softwareccs: Software and its engineering Interactive games
1. Introduction
In the maintenance phase of the game development cycle, it is crucial to provide updates that improve the game to maintain player retention and enhance engagement (Zhong_Xu_2021, ). These game updates should also fix bugs for improved user experience. As a type of bug, crash bugs are especially severe because they prevent the user from experiencing the game. They are also difficult to fix, being the most recurring bug type even after updates addressing the bug (Truelove_Santana_De_Almeida_Ahmed_2021, ).
A primary source of information for these updates is the user-submitted bug reports. A prime example is Mojira [^1] https://bugs.mojang.com/issues, the public bug tracker for Minecraft developer Mojang Studios. Mojira hosts more than 450,000 bug reports, with over half related to Minecraft: Java Edition [^2] https://www.minecraft.net/en-us/about-minecraft. However, many of these reports are of low quality, often missing essential components like steps to reproduce (S2R), expected outcomes, and can have mistakes in them (Zimmermann_Premraj_Bettenburg_Just_Schroter_Weiss_2010, ; Soltani_Hermans_Bäck_2020, ). The studies showing low-quality bug reports across numerous projects suggest that Mojira could potentially face similar challenges.
Due to the excessive amount of labor involved, game testing is an entire field in itself, and it is a notoriously manual process that takes a significant amount of time (Politowski_Petrillo_Gueheneuc_2021, ). Most games lack automated test frameworks because implementing and generalizing such systems is challenging. For example, Minecraft’s internal unit test framework GameTest [^3] https://www.minecraft.net/en-us/creator/article/get-started-gametest-framework lacks the tools to reproduce UI movement and precise player interactions, only covering a part of what is a huge sandbox game.
Looking at a more general usage, automated bug reproduction systems have been developed (Kang_Yoon_Yoo_2023, ). With the advent of large language models (LLMs) that are capable of tool calling (i.e., interacting with external tools/APIs) (qin2024toollearningfoundationmodels, ) and code generation (jiang2024surveylargelanguagemodels, ), the implementations of these systems have become more general. Similar works (Wang_Zhao_Feng_Zhang_Halfond, ) exist for reproducing bugs in Android apps through interacting with the available GUI. These works are limited in interacting with the game, which needs action taking in a complex environment. To our knowledge, there are no works for automating bug reproduction in an open-ended game environment like Minecraft.
In this paper, we present BugCraft, an agentic bug reproduction framework for Minecraft. BugCraft can take in the initial bug report, synthesize high-quality steps to reproduce bugs alongside relevant game data from external sources and reproduce them step by step through scripting. In the end, our framework generates a log that can reproduce the bug. In addition, due to the lack of a dataset of crash bugs in a game, we present BugCraft-Bench, a crash bug dataset collected through Mojira. Our dataset contains 86 manually confirmed, reproducible crash bug reports. Finally, our reproduction package, alongside with our results can be found at: https://figshare.com/s/70b5d141e69ff822cd40.
Figure 1. The BugCraft framework, illustrating the two-stage process of S2R synthesis and action model execution.
Our paper begins by discussing related work in Section 2, followed by an overview of the framework in Section 3. We define our research questions in Section 3.3, evaluate the framework in Section 4, and discuss our findings in Section 5. The paper mentions the threats to validity in Section 6, and concludes with a summary of contributions in Section 7.
2. Related Work
This section provides an overview of related topics, focusing on bug reproduction in games and the development of agents powered by LLMs. We also discuss similar research efforts relevant to our framework.
2.1. Bug Reproduction in Games
Game testing is traditionally a manual and time-consuming process, and the primary approach involves playtesting, in which testers manually play the game to identify issues (Politowski_Petrillo_Gueheneuc_2021, ). Automated unit testing is often challenging for game developers due to the dynamic nature of game code and the complexity involved in capturing game states programmatically (Pascarella_Palomba_Di_Penta_Bacchelli_2018, ). In continuous development cycles, it becomes essential for developers to address existing bugs through regular updates to improve the gameplay experience. However, certain types of bugs, particularly crash bugs, are prone to reoccurrence even after an update is released to fix them (Truelove_Santana_De_Almeida_Ahmed_2021, ). One developer in a recent survey attributed this issue to limited or non-available testing time, highlighting the constraints of manual testing in the fast-paced, deadline-driven environment typical of game development.
Wuji (Zheng_Xie_Su_Ma_Hao_Meng_Liu_Shen_Chen_Fan_2019, ) represents the closest related work in automated game testing, using deep reinforcement learning to explore the game model to discover new bugs. However, Wuji’s capabilities are limited to predefined game areas and narrow action spaces. It requires manual environment setup without the ability to follow specific test prompts. Furthermore, Wuji operates in the domain of bug finding, while our framework focuses on reproducing known bugs. Our approach provides end-to-end bug reproduction capabilities, starting from game initialization and covering specific scenarios, enabling a more comprehensive approach to addressing reported issues.
2.2. Automated Bug Reproduction Using LLMs
Automated bug reproduction entails converting unstructured bug reports into executable scripts reliably replicating the reported issues. LLMs have shown considerable success in generating code, particularly in creating unit tests (Kang_Yoon_Yoo_2023, ) and enhancing bug report details (Bo_Ji_Sun_Zhang_Wu_Wei_2024, ). Recent frameworks leveraging LLMs have been developed to automate bug reproduction in mobile applications, particularly within Android environments, by interacting with the GUI (Wang_Zhao_Feng_Zhang_Halfond, ). These frameworks enable a series of structured actions where the LLM makes decisions based on the bug report, GUI interactions, and prior actions while additional contextual information is provided as needed by the framework. Frameworks in (Wang_Zhao_Feng_Zhang_Halfond, ) and (Feng_Chen_2024, ) rely on methods that convert the GUI into textual representations, allowing the LLM to interact with and manipulate the interface through text-based commands. Additionally, a study (Liu_Li_Chen_Wang_Wu_Wang_Hu_Wang_2024, ) incorporates visual inputs, enhancing the LLM’s capacity to interpret and interact with complex GUIs. This approach has achieved performance comparable to other vision-based methods, highlighting the potential for LLMs to effectively combine textual and visual data in automated bug reproduction tasks.
2.3. LLM Agents
LLMs have shown great success at acting as artificial agents, which are entities capable of perceiving their surroundings, making decisions, and taking actions in response (xi2023rise, ; yao2023react, ). The agent framework involves equipping LLMs with tools, often in the form of APIs, to expand their action space and enable them to interact effectively with the environment. The most important part is that the LLMs were able to do this mainly by utilizing in-context learning (tombrownfewshot, ), alleviating the need to train specialized models for each in-domain task. Moreover, LLMs have also shown to be successful at taking action in an open-ended environment. This ability to act extends even to games like Minecraft, where the Voyager framework leveraged an LLM agent to achieve in-game progress milestones significantly faster than previous methods (wang2023voyager, ).
3. Methodology
Our experiment utilizes a custom dataset of Minecraft crash bug reports, processed through a custom framework called BugCraft, which operates in two main phases: creating the Steps to Reproduce (S2R) and executing the S2R (step clusters). Figure 3 provides a high-level overview of our framework. The following sections will elaborate on its components: we begin with the dataset creation process (subsection 3.1), followed by a brief introduction to Bugcraft (subsection 3.2). We then describe the preprocessing of a bug report (subsection 3.2.1), the Step Synthesizer component (subsection 3.2.2), the Action Model component (subsection 3.2.3), the general implementation details of the entire framework (subsection 3.2.4) and finally the evaluation setup (subsection 3.3).
3.1. Data Curation
To our knowledge, no readily available dataset of Minecraft bug reports exists. Therefore, to quantitatively measure the efficiency of our framework, we constructed our own dataset of crash bug reports. Furthermore, we gathered data from the Minecraft Wiki to enrich our framework with in-game information.
3.1.1. BugCraft-Bench Dataset Curation
We built a custom dataset consisting of 86 Minecraft: Java Edition crash bug reports to evaluate our framework. These reports were collected from Mojira, Minecraft’s official bug tracker, utilizing Jira’s REST API. We employed a finegrained query to retrieve all reports classified under the ”Crash” category. Moreover, for each report, we gathered all associated comments, attachments, and external links (e.g., YouTube links), which may contain relevant files or videos. These were identified using regular expressions (Regex).
Figure 2. Bug Report Codenamed MC-276621
In Figure 2, we present a sample bug report, identified as MC-276621444https://bugs.mojang.com/browse/MC-276621, which was verified and resolved by Mojang in a subsequent snapshot. This report illustrates a bug that can be reproduced through a specific combination of commands and UI interactions. Additionally, another user, contributed steps-to-reproduce (S2R), which were initially missing from the report. Importantly, the BR in its final state lacks commands that would make reproducing the bug easier. Furthermore, though not relevant to this particular report, some bug reports, such as MC-275657555https://bugs.mojang.com/browse/MC-275657, contain an implicit assumption that a version-specific setting is enabled without explicitly stating it, creating ambiguity during reproduction.
Each report in the dataset was categorized according to its ”Confirmation Status” label in Mojira, which includes four possible states: ”Confirmed,” ”Community Consensus,” ”Plausible,” and ”Unconfirmed.” We primarily focused on ”Confirmed” and ”Community Consensus” reports, as these have been verified by the game’s developer, Mojang Studios, making them reliable ground truth instances. To focus solely on singleplayer bugs, we excluded reports related to servers, multiplayer interactions, and those requiring external files loaded into the game (such as datapacks, worlds, or world setting JSONs).
To ensure the quality and relevance of our dataset, we employed a multi-stage filtration process, focusing specifically on single-player crash bugs. We began by targeting the most recent 300 (limited the size of the dataset because of resource constraints) crash bug reports from Mojira’s database following our query results, spanning from the most recent, MC-277967666https://bugs.mojang.com/browse/MC-277967 (resolved on 05/Nov/24, affecting version 24w44a), to MC-145102777https://bugs.mojang.com/browse/MC-145102 (created on 19/Feb/19, affecting version 19w08b), covering a total of 70 different game versions. We then downloaded all associated data for each of these 300 reports, including comments, attachments, and external links. Next, we undertook a manual review, a crucial step that involved categorizing each report based on several criteria. Our focus was to isolate single-player crash bugs, so we excluded reports involving multiplayer aspects like servers or Realms (a subscription based server service) [^8] https://www.minecraft.net/en-us/realms, those requiring interactions between different game versions, and any that depended on external resources such as datapacks, custom worlds, resource packs, or specific world settings. We also removed reports related to issues outside the game itself or those that were machine-dependent. Furthermore, reports were discarded if they were considered low quality due to inadequate information or if the final version differed so much from the initial draft that it could be considered a new submission. This meticulous review resulted in a refined set of 141 reports. Due to resource constraints and the significant cost associated with the LLM API calls, we chose to evaluate only a portion of the full dataset. As such, from the initial pool of 141 reports, we randomly selected a subset of 100 to create our final dataset. This dataset was used for detailed analysis and the evaluation of our framework.
3.1.2. Minecraft Wiki Dataset
We use the official Minecraft Wiki [^9] https://minecraft.wiki site as our knowledge base. Since the base of the site used MediaWiki, we are able to get the wiki’s text using the provided wiki API. We pulled every page in the wiki that was not a redirect page. We stored each page as a separate file and used fuzzy matching with the page titles for retrieval. Using the wiki as our knowledge base allows us to circumvent reliance on our underlying LLM’s pre-training data.
3.2. BugCraft Framework
Figure 3. BugCraft Framework
The BugCraft framework (Figure 3) is designed for automated bug reproduction and comprises two core components: the Step Synthesizer and the Action Model. These components work in tandem, with the Step Synthesizer generating a sequence of steps and the Action Model providing the executable actions, to transform raw, unstructured bug report data into executable scripts for accurate bug replication. We outline each component below.
3.2.1. Preprocessing
The initial stages of our process involve a crucial preprocessing phase applied to the raw bug report data. This phase begins with extracting the entirety of the textual content, encompassing the very first submission of the bug report and all subsequent comments. It’s important to understand that we specifically focus on these initial versions because they represent the most challenging scenario for our framework. Specifically, we analyze the state of the bug report as it was initially submitted by the user before any modifications or refinements were made by the Mojira team. These initial reports are often less structured, potentially containing incomplete information or ambiguous language, reflecting the user’s initial understanding and description of the issue. This contrasts significantly with later, revised versions, which benefit from the input and clarifications provided by experienced team members. The core challenge, and our framework’s primary objective, is accurately interpreting and addressing these raw, initial reports. Subsequent refinements often simplify the report, making the issue easier to understand, but our focus is on the initial state of the report created by the user.
While the Mojira API lacks a direct method for retrieving previous versions, it does provide access to the issue’s changelog, detailing the history of each field. We developed a back-tracker script to parse these fields and reconstruct the bug report at a specific point in its changelog. Comments are not included in the changelog, so we delete any comments that were added after the selected reconstruction step. To maximize the availability of information, the back-tracker stops at the latest state of the BR that has not been touched by a Mojira team member or a trusted individual (Users with [Helper] role). Using Mojira’s role-based system, we specifically target the last version prior to any changes made by users with the roles [Mod], [Helper], or [Mojang].
3.2.2. Step Synthesizer
Figure 4. Step Synthesizer Component
The Step Synthesizer component seen in Figure 4 is responsible for converting the unstructured text of bug reports into high-quality S2R steps, augmented with relevant external information. For the entire step synthesizing process, we use the same MLLM (Multimodal Large Language Model), which we call the Step LLM. Further details regarding the prompts utilized in the Step Synthesizer are provided in the reproduction package. This process occurs in the following stages:
Knowledge Augmentation: Prior to generating the S2R steps, we aim to identify relevant wiki pages. This process begins by prompting the Step LLM with the bug report to perform named entity recognition, resulting in a list of entities. Fuzzy matching is then employed to compare this list against wiki page titles, yielding a set of potentially relevant wiki page titles. Then, the Step LLM selects the most relevant titles from the set. In addition to these specifically identified pages, we incorporate generally relevant pages, such as those pertaining to the bug report’s version, general game mechanics, and commands. Once all titles are selected, the corresponding pages are fetched, and individual analysis trajectories are generated for each page. Inspired by the self-reasoning paper (xia2024improvingretrievalaugmentedlanguage, ), which generates a reasoning trajectory for each retrieved chunk, we adopt a similar approach to analyze each page. An analysis trajectory is a step-by-step analysis generated by the Step LLM that examines the relevance of a specific external information source (e.g., a wiki page) to the bug report, highlighting key details and their potential implications for bug reproduction. This is crucial for handling the potential length of the pages and ensuring that only relevant information is cited. The LLM is also instructed to disregard any page deemed irrelevant. The resulting analysis trajectories, representing the relevant information extracted from the wiki pages, are appended to the end of every LLM call unless otherwise specified.
Initial S2R Generation: Leveraging the bug report text, comments, and curated web content, we generate an initial S2R prompting the Step LLM. This S2R is crafted to address the most critical aspects of bug reproduction while simultaneously synthesizing the unstructured text from the bug report and comments into a single, structured list. Furthermore, the use of Step LLM during this generation process enhances the S2R’s effectiveness by enabling the creation of more readily executable steps by eliminating ambiguity and promoting the use of easier and more reliable alternate steps compared to the original bug report.
S2R Enhancement: Following the initial one-shot generation of the S2R, we subject it to a two-stage critique process, each stage utilizing a separate Step LLM call. This critique specifically targets inconsistencies within the plan and the predicted behaviors of in-game entities (mobs) [^10] https://minecraft.wiki/w/Mob. The initial S2R may contain errors regarding common mob behaviors, such as targeting, movement, and other related factors. To address this, we employ few-shot (tombrownfewshot, ) examples to guide the Step LLM in generating suggestions for improvement. Subsequently, we utilize two additional Step LLM calls to rewrite the S2R, placing emphasis on incorporating the generated suggestions. This iterative refinement process (madaan2023selfrefineiterativerefinementselffeedback, ) ensures the final S2R is comprehensive, internally consistent, and accurately reflects the expected behaviors of other agents in the game, such as mobs.
Step Clustering: The finalized S2R is then used as input for a Step LLM prompt that instructs it to cluster the steps. This process empirically yields 2-4 clusters, each with a descriptive title and a list of its constituent steps. This clustering improves the readability and digestibility of the S2R. Following cluster generation, a critique is performed to identify any misplaced steps. Based on this critique, steps are reordered as needed. The outcome of this process is a refined and finalized S2R, optimized for clarity and downstream tasks. For instance, in MC-276621, our system effectively integrated the initial bug report and subsequent comment, enhancing the steps to reproduce with relevant commands for easier replication (Figure 5).
Figure 5. Initial MC-276621 Report and Our Output
Upon completion, the Step Synthesizer yields a refined, fully annotated S2R set, ready for execution by the Action Model.
3.2.3. Action Model
Figure 6. Action Model Component
The Action Model component, illustrated in Figure 6, is responsible for translating the refined S2R into executable scripts. These scripts interact with the game interface, encompassing both the main menu and in-game elements. To achieve this, we developed a custom macro API, as existing solutions lacked a unified approach for handling both GUI and in-game interactions. Furthermore, the development of a custom macro API enabled the inclusion of game-specific methods tailored for frequently used interactions. To enhance the LLM’s ability to identify and interact with specific GUI elements, we integrated OmniParser (lu2024omniparserpurevisionbased, ). OmniParser is, as of November 2024, a state-of-the-art model framework for parsing UI screenshots into structured elements. It excels at reliably identifying interactable icons and texts within the user interface and understanding the semantics of various elements in a screenshot while providing annotations.
3.2.3.1 Base Macro API
The custom API uses a JSON-based REST schema, allowing the Action Model to interact with the game environment. The API supports the following methods:
press(key: Key—List[Key], time: Optional[float]): Handles keyboard and mouse input. If given multiple keys, it presses them simultaneously. The optional time parameter dictates the duration (in seconds) for which the keys are held down; if omitted, the keys are pressed and immediately released.
write(str: string): Types the provided string into the currently focused text box using the keyboard.
command(instruction: String): Executes Minecraft commands. Commands are restricted to non-menu contexts. This is the shortcut for the LLM to open the chat, write the command, and then send the command message.
click(coordinates: Point): Moves the mouse and clicks at the specified coordinate.
Each JSON request can encapsulate multiple methods that are executed sequentially. This allows for the orchestration of complex actions that involve menu navigation, gameplay commands, and GUI interactions. However, the click method is not directly used by the action-taking LLM. Instead, a click_place method is provided. This method acts as an abstraction layer, bridging the OmniParser annotations with the API. The LLM specifies a particular bounding box identified by OmniParser, and the click_place method translates this into a click command, using the center coordinates of the specified bounding box.
3.2.3.2 Action Model Framework
Prior to initiating the logic loop, the game is launched using the version specified in the bug report, and the macro API is run in a separate thread. Finally, OmniParser is initialized and made ready to annotate images.
Example Trajectory Iteration Current Step Title: Setup Minecraft Environment Thought: We have accessed the ’More’ options menu. The next step is to set the world type to ’Flatworld’. We need to click on ’World’ to access the world type settings. The correct index for ’World’ is 12. Action: Clicked the place that had content: world. at coordinates: [x1: 0.41, y1: 0.06, x2: 0.60, y2: 0.12] Reflection: The action was successful. The screen now shows the ’World’ options menu, which includes the ’World Type’ setting. This indicates that the ’World’ button was clicked correctly, allowing access to the world type settings. This matches the expected outcome, as the next step is to set the world type to ’Flatworld’. Classification: SUCCESS
For the Action Model loop, we implement a logic similar to ReAct (yao2023react, ). Further details regarding the prompts utilized in the Action Model are provided in the reproduction package. For the entire action-taking process, we use the same MLLM, which we call the Action LLM. To execute each S2R, the Action Model follows this loop:
OmniParser Annotation: Before each action, a screenshot of the Minecraft window is captured for annotation by OmniParser. OmniParser analyzes the image, identifying all icons and text elements. Additionally, it classifies each element as either interactable or non-interactable. The resulting data is then formatted into a structured markdown table.
Thought & Action Generation: Prior to executing an action, the Action LLM is prompted to generate both a ”thought,” outlining its intended course of action, and the corresponding action itself in a single step. This prompt is supplied with the current trajectory (comprising previous thoughts, actions, and feedback), the annotated table generated by OmniParser, the corresponding game screenshot, and relevant examples. The combination of visual input from the game image and the structured information from the annotation table facilitates the Action LLM’s ability to identify and interact with specific game elements.
Action Verification: Following the generation of the thought and action, the Action LLM is prompted to verify whether the proposed thought and action are correct (madaan2023selfrefineiterativerefinementselffeedback, ) and conducive to progressing the game state. This verification step acts as a crucial safeguard against errors and infinite loops, where the agent might repeatedly execute the same ineffective action. If the verification fails, the generated feedback, indicating the reason for failure, is used to prompt the generation of a new thought and action pair. This revised pair is then passed to the Macro API for execution.
Feedback Generation: After each action, a new screenshot is taken. The Action LLM then assesses the outcome, considering the prior thought and action. This involves judging if the thought was achieved and if the current step cluster is nearer completion, with a rationale. If all steps in the cluster are deemed complete, the Action LLM can signal a move to the next. This triggers a second Action LLM call to independently verify completion as a redundant check. If verification fails, the active cluster stays the same.
The action model loop runs until either the game crashes or a preset iteration limit is reached. Because the macro API logs each action, it generates a log capable of consistently reproducing the bug, assuming the bug isn’t dependent on randomness.
3.2.4. S2R & Action Model Implementation Details
We implement the framework using Python. We use LangChain [^11] https://www.langchain.com/, to implement the LLM calling logic. For all LLM calls, we keep the temperature at 0 to maintain maximum determinism. For the Action Model, we only keep the last 25 iterations in Action LLM’s context to prevent context length overflow and set the iteration limit as 30. Moreover, while formatting the trajectory iterations, instead of putting the commands, we put verbal representations of what that command means. We do this to increase the interpretability of the generated actions for the LLM.
3.3. Evaluation Setup
To evaluate BugCraft, we pose two research questions about its components, considering real-world usage:
RQ1: How effective is our step synthesizer for producing correct plans?
RQ2: How effective is BugCraft at reproducing crash-inducing bug reports in an end-to-end manner?
To address RQ1 and RQ2, we conducted experiments using a curated subset of our crash bug dataset. This subset comprises 100 bug reports, carefully selected to represent a variety of scenarios that require interaction with both GUI and in-game world elements, as mentioned in section 3.1.1. The selection of this refined dataset was driven by the resource-intensive nature of running full evaluations, making it impractical to test all selected bug reports within the scope of this study.
3.3.1. Evaluation of the Step Synthesizer
To evaluate the accuracy of the step clusters generated by the Step Synthesizer, we manually reviewed each S2R. This involved attempting to reproduce the bug in Minecraft by strictly following the generated steps. If ambiguities or obstacles were encountered, we first consulted the original bug report for clarification. If the original report was insufficient, we referred to the final version of the bug report on Mojira and other relevant online resources (e.g., Minecraft Wiki) to understand the intended reproduction process.
Each step cluster was then labeled as either ”True”, ”Faulty”, or ”Irreproducible”.
  • True: Following the step cluster leads to successful bug reproduction. Minor deviations or extra guidance steps for the LLMs are permitted.
  • Faulty: The step cluster is flawed and prevents bug reproduction. It may be missing crucial steps, contain incorrect commands, or have logical inconsistencies.
  • Irreproducible: The bug cannot be reproduced due to reasons unrelated to the Step Synthesizer’s output. Examples include dependence on outdated versions, specific hardware, or fundamental impossibility.
To ensure the reliability of our labeling, particularly the distinction between ”Faulty” and ”Irreproducible,” we conducted an inter-rater agreement analysis. Prior to the main labeling task, we jointly inspected the results of a test run to establish the labeling categories and refine our understanding of the criteria. Subsequently, we independently labeled the 100 bug reports in our run for the step synthesizer. Inter-rater agreement was measured using both percentage agreement and Cohen’s Kappa (McHugh_2012cohen, ). The results of this analysis, including a confusion matrix, are presented 4.2.1.
Faulty step clusters were further analyzed to identify the specific type of error. Each faulty step cluster was assigned at least one of the following error categories:
  1. (1)
    Wrong Command: An S2R is categorized here if it includes commands that are not recognized or parseable by Minecraft. This often occurs during the S2R enhancement phase, where the LLM may rely on its pre-trained data, which might not encompass newer command syntax, resulting in the substitution of older, invalid command formats.
  2. (2)
    Missing Step: An S2R falls under this category when it lacks a critical step, rendering the subsequent steps implausible. This issue was more prevalent with bug reports from newer Minecraft versions. This could be attributed to the increasing complexity of the game with each update and the training data cutoff date of GPT-4o (October 2023)(openai2024gpt4ocard, ), potentially leading to an incomplete understanding of the latest game mechanics.
  3. (3)
    Logic Error: This category includes steps that are impossible, significantly hinder reproduction, or are internally contradictory within the plan itself For example, simultaneously setting the time to day and spawning hostile mobs (which would burn in daylight), or attempting actions restricted by an open inventory screen.
To ensure the reliability of our labeling, we conducted an inter-rater agreement analysis. Prior to the main labeling task, we jointly inspected the results of a test run to establish the labeling categories and refine our understanding of the criteria. Subsequently, we independently labeled the 100 bug reports in our main dataset. Inter-rater agreement was measured using both percentage agreement and Cohen’s Kappa. The results of this analysis are presented in  4.2.
3.3.2. Evaluation of the Action Model
To evaluate the Action Model’s ability to reproduce bugs, we conducted a comprehensive analysis of its performance on the subset of 86 bug reports that were deemed reproducible during the Step Synthesizer evaluation (Section 3.3.1). Each reproducible report was processed end-to-end, and the Action Model’s success or failure in reproducing the bug was recorded.
3.3.3. Determining Action Model Success or Failure
The initial assessment of the Action Model’s performance focused on whether it successfully reproduced the bug or not. We classified success as instances where the game crashed as described. Failures occurred when no crash was observed. Due to the straightforward nature of this assessment, a formal inter-rater agreement was not deemed necessary.
3.3.4. Failure Analysis and Inter-Rater Agreement
To understand the reasons behind failures, we performed a detailed examination of each unsuccessful attempt. This involved identifying recurring patterns and common failure modes, which informed the development of a structured categorization rubric. Just like we did with the Step Synthesizer rubric, we collaboratively analyzed the results of the same test run to define the failure categories, refine our understanding of the criteria, and establish a shared understanding of potential edge cases in applying the rubric. We also used our understanding of common agent mistakes to enhance the rubric. The results of this analysis, are presented in Section 4.3.
3.3.4.1 Action Model Grading Rubric
A. On Step Synthesizer Success
  1. (1)
    Success: The agent successfully reproduces the bug by following the provided plan.
  2. (2)
    Failure: The agent fails to reproduce the bug.
    1. (a)
      Agent Incapability: The agent fails to reproduce the issue due to its own limitations, despite a correct plan and sufficient framework capabilities.
      1. (i)
        Stuck in Loop: The agent enters a repetitive loop, hindering progress.
        1. (A)
          Menu Loop: The agent is trapped within the game’s menu system, particularly the ”Create World” screen in newer versions.
        2. (B)
          Command Loop: The agent repeatedly attempts to execute the same incorrect or unparseable command.
        3. (C)
          Death Loop: The agent continuously dies due to in-game factors, such as hostile mobs, preventing it from acting.
      2. (ii)
        Poor Decision Making: The agent’s suboptimal choices lead to failure. This includes neglecting environmental factors, skipping crucial steps, or taking misguided actions, as detailed in the initial category descriptions.
    2. (b)
      Framework Incapability: The framework’s limitations make bug reproduction impossible or extremely difficult. This includes the inability to execute actions within tight time constraints or to interact with specific menu elements.
  3. (3)
    Error: The action model encounters an internal error and crashes (e.g., the LLM failing to follow the JSON schema).
B. On Step Synthesizer Failure
  1. (1)
    Recovery: Despite a faulty plan from the Step Synthesizer, the agent successfully reproduces the bug.
  2. (2)
    Total Failure: Both the Step Synthesizer and the Action Model fail to reproduce the bug.
  3. (3)
    Error: The action model encounters an internal error and crashes (e.g., Pydantic error).
4. Results
This section presents the results of our evaluation of BugCraft, focusing on the two main components: the Step Synthesizer and the Action Model. We first analyze the Step Synthesizer’s ability to generate accurate bug reproduction plans, followed by an examination of the Action Model’s performance in executing these plans within the complete end-to-end pipeline.
4.1. Experimental Setup
Our experiments were conducted on a high-performance machine running Windows 11, equipped with 32GB of RAM, an Intel i9-14900HX CPU, and an NVIDIA RTX 4090 Mobile GPU.
We used OpenAI’s GPT-4o (specifically, the gpt-4o-2024-11-20 version) as the underlying language model for both the Step Synthesizer and Action Model components. We chose this model due to its state-of-the-art performance and widespread adoption (openai2024gpt4ocard, ). The temperature parameter for all LLM calls was set to 0 to ensure deterministic and consistent outputs(renze2024effectsamplingtemperatureproblem, ).
The complete testing process for the 100 bug reports took approximately 36 hours. This duration is attributed to the inherently sequential nature of testing each bug report, as parallelization was not feasible. Individual runs could take up to 30 minutes, depending on the complexity of the bug and the actions required for reproduction. [^12] https://minecraft.wiki/w/Java_Edition_19w13b
4.2. Step Synthesizer Analysis
Table 1. Step Synthesizer Planning Correction Analysis
Outcome Count Percentage (%)
True 57 66.28%
Faulty 29 33.72%
Faulty Analysis Breakdown*:
        Wrong Command 14 48.28%
        Missing Step 10 34.49%
        Logic Error 9 31.03%
Irreproducible 14 -
*One faulty case may belong to multiple categories.
Table 1 summarizes the results of our Step Synthesizer evaluation, conducted according to the methodology outlined in Section 3.3.1. As detailed in our evaluation setup, we manually tested the generated step clusters by attempting to reproduce the bugs in the corresponding Minecraft game versions. The step clusters were categorized as ”True” if they could successfully lead to bug reproduction, ”Faulty” if they contained errors preventing reproduction, or ”Irreproducible” if the bug itself could not be reproduced even by using the initial BR, which is not a fault of the step synthesizer.
The key findings, as shown in the table, are that 66.28% of the generated step clusters were deemed ”True”, indicating a reasonable level of accuracy in the Step Synthesizer’s plan generation. However, 33.72% of the plans were categorized as ”Faulty.” Further analysis of these faulty plans revealed that the most common error was ”Wrong Command” (48.28%), followed by ”Missing Step” (34.49%) and ”Logic Error” (31.03%). Moreover, we classified 14 bug reports as ”Irreproducible,” due to factors such as dependence on specific hardware configurations or errors in the original bug report, underscoring the complexities of bug reproduction.
4.2.1. Inter-Rater Agreement for Step Synthesizer
To ensure the reliability of our manual labeling of step clusters, we conducted an inter-rater agreement analysis. The results are presented in Figure 7. The confusion matrix shows the agreement and disagreement between the two authors on the classification of step clusters as ”True,” ”Faulty,” or ”Irreproducible.” A Cohen’s Kappa of 0.70 and a percentage agreement of 83.0% were achieved, indicating substantial agreement.
Figure 7. Confusion matrix of bug report reproducibility showing inter-rater agreement
4.3. Action Model Analysis
Table 2. Experimental Results and Analysis of Failure Cases
Category Count Percentage (%)
A. Step Synthesizer Success 57 100.0%
   Success (Bug Reproduction) 22 38.60%
   Failure 32 56.14%
   Breakdown of Failures:
   $\circ$ Agent Incapability 29 50.88%
    - Stuck in Loop 11 19.30%
     $\bullet$ Menu Loop 8 14.04%
     $\bullet$ Command Loop 2 3.51%
     $\bullet$ Death Loop 1 1.75%
    - Poor Decision Making 17 29.82%
   $\circ$ Framework Incapability 4 7.02%
   Error 3 5.26%
B. Step Synthesizer Failure 29 100.0%
   Recovery 4 13.79%
   Failure 24 82.76%
   Error 1 3.45%
Overall Totals
   Total Trials 86 100.0%
   Total Success 26 30.23%
   Total Failure 56 65.12%
   Total Errors* 4 4.65%
    Three errors were from internal crashes (LLM failed to give the output in correct JSON schema) and one error was from the technical issues that the potential bug could cause.
Table 2 summarizes the results of our Action Model’s performance on 86 bug reports, conducted as per the methodology in Section 3.3.4.1. Initially, 100 bug reports were considered. However, 14 were deemed ”Irreproducible” during the Step Synthesizer evaluation (see Section 4.2), and an additional 4 were excluded due to errors encountered during the Action Model runs: 3 internal framework crashes and 1 technical issue the potential bug could cause to our setup. These errors prevented the runs from concluding successfully. Therefore, the Action Model ultimately attempted to reproduce 86 bug reports, including those with faulty step clusters identified in the Step Synthesizer analysis. As explained in our evaluation setup (Section 3.3.2), we manually categorized the Action Model’s failures and also specifically inspected the bug reports whose step clusters were generated as ”Faulty”. In these cases, there was a chance for our agent to explore and find the correct way to reproduce the bug by utilizing its inherent knowledge, and the potential feedback it could get from the game.
In our comprehensive end-to-end testing, the Action Model successfully reproduced 26 out of 86 bugs (30.23%). When provided with correct step clusters (”Step Synthesizer Success” cases), the Action Model achieved a 38.60% success rate (22 out of 57). However, when the step clusters were faulty (”Step Synthesizer Failure” cases), the Action Model only managed to recover and reproduce the bug in 4 out of 29 cases (13.79%). These recovery cases are particularly significant because they demonstrate the agent’s ability to overcome errors in the provided plan and successfully reproduce the bug using other available information.
Analyzing the failures when correct step clusters were provided, we found that ”Agent Incapability” was the most common cause (50.88%), with the failures being distributed between ”Poor Decision Making” (29.82%) and ”Stuck in Loop” (19.30%). ”Framework Incapability” accounted for a smaller portion of the failures (7.02%). In cases where the Step Synthesizer failed to generate correct steps, and the Action Model also failed to recover, the vast majority (82.76%) were attributed to the faulty steps.
5. Discussion
This study aimed to evaluate the BugCraft framework through two primary research questions, focusing on the performance of its two main components: the Step Synthesizer and the Action Model. Our findings provide valuable insights into the capabilities and limitations of using LLMs for automated bug reproduction in complex game environments like Minecraft.
5.1. Addressing RQ1
The Step Synthesizer demonstrated a promising ability to transform raw, unstructured bug report data into structured, actionable steps for bug reproduction. It achieved a 66.28% success rate in generating accurate plans (step clusters), as evaluated through the methodology outlined in Section 3.3.1. This indicates that the component is generally effective in understanding the reported issue, extracting relevant information, and synthesizing it into a coherent sequence of actions. The strong correlation between successful plan generation and the Action Model’s success further validates the Step Synthesizer’s performance, highlighting its crucial role in the overall bug reproduction pipeline.
However, the 33.72% failure rate reveals that challenges remain. The analysis of faulty plans identified three main error categories: ”Wrong Command” (48.28%), ”Missing Step” (34.49%), and ”Logic Error” (31.03%). ”Wrong Command” errors often stemmed from the LLM relying on outdated command syntax from its pre-training data, especially for newer Minecraft versions. This was particularly evident in the LLM’s struggle to generate correct item tags, which have a detailed schema that changes frequently between versions [^13] https://minecraft.wiki/w/Data_component_format. Additionally, some hallucinated commands were generated that never existed in the game. ”Missing Step” errors indicated that the Step Synthesizer sometimes failed to capture all the crucial details needed for successful reproduction, particularly for complex bugs or those involving recent game mechanics. This often manifested as the LLM failing to adequately track and account for the evolving game state. ”Logic Error” cases demonstrated challenges in ensuring the generated plan’s internal consistency and adherence to game rules.
The Step Synthesizer offers a good foundation for bug reproduction planning, but needs improvement. It requires better integration of version-specific game information, command verification (e.g., syntax parsing), and logical consistency checks. Solutions could include enhancing external data integration and incorporating game environment feedback to refine plans dynamically.
5.2. Addressing RQ2
BugCraft achieved a 30.23% success rate in end-to-end reproduction of crash-inducing bug reports from our curated dataset. This demonstrates a tangible capability for automated game bug reproduction, showcasing the potential of LLMs to bridge the gap between natural language descriptions of bugs and concrete actions within a complex game environment. While this success rate indicates the feasibility of our approach, the corresponding 65.12% failure rate highlights the inherent challenges in automating bug reproduction within such environments.
The primary bottleneck in our end-to-end pipeline was the Action Model. Even when provided with accurate step-by-step plans from the Step Synthesizer, the Action Model still failed to reproduce the bug in a majority of cases (56.14%). This underscores the difficulty of translating even well-defined plans into successful in-game actions. Agent incapability was the most significant factor hindering the Action Model’s performance. This category encompassed both ”Poor Decision Making” (29.82%) and being ”Stuck in Loops” (19.30%). ”Poor Decision Making” highlighted the challenges faced by the Action LLM in navigating the complex, visually-rich Minecraft environment. Operating solely on visual input, without a deeper symbolic understanding of the game state, the LLM sometimes struggled to make optimal choices, overlooking environmental cues or deviating from the intended plan. The ”Stuck in Loop” failures further illustrated these challenges, manifesting as ”Menu Loops” (14.04%), ”Command Loops” (3.51%), and even a ”Death Loop” (1.75%). ”Menu Loops,” particularly within the ”Create World” screen, indicated difficulties in reliably navigating complex GUI structures. ”Command Loops” arose from issues in command parsing or execution, with the agent repeatedly attempting unparseable or ineffective commands. The single ”Death Loop” instance was caused by hostile mobs continuously killing the player, preventing progress.
Beyond agent-specific limitations, ”Framework Incapability” contributed to 7.02% of failures when correct steps were available. This suggests that while our custom API provided a foundational level of interaction, it lacked the granularity or robustness needed for certain bug reproduction scenarios. Limitations in precise mouse control, rapid action execution, or interaction with specific UI elements within the game hindered the Action Model’s ability.
Notably, the Action Model successfully reproduced bugs in 13.79% of cases despite faulty plans from the Step Synthesizer. These ”Recovery” cases suggest the agent can leverage its knowledge and context to overcome some inaccuracies, hinting at the potential for a more robust and adaptable Action Model capable of problem-solving. However, the low recovery rate underscores the continued importance of accurate Step Synthesizer plans.
While BugCraft shows promise for automated game bug reproduction, improvements in agent capabilities and framework robustness are essential. Future work should focus on enhancing the agent’s understanding of the game environment, possibly via symbolic or hybrid approaches. Improving the game interaction API for better control/feedback, and adding more robust error detection and recovery, could also boost performance.
6. Threats to Validity
6.1. Construct Validity
A potential threat to construct validity stems from using the Minecraft Wiki for knowledge augmentation. While BugCraft might generalize to other games, especially those with developer support, the exceptionally detailed Minecraft Wiki may not be representative of typical game documentation. This, combined with Minecraft’s popularity, likely led to substantial Minecraft data in the LLM’s pre-training. This pre-existing knowledge could inflate the LLM’s understanding of Minecraft, overestimating BugCraft’s performance. Games with less documentation or popularity might see reduced LLM effectiveness, limiting generalizability. Furthermore, the LLMs might have been trained on the final, corrected versions of Mojira bug reports, even though we used initial, unedited versions, potentially overestimating the LLM’s ability to handle raw bug reports.
6.2. Internal Validity
The internal validity of our study is potentially affected by the inherent stochasticity of LLMs, which typically produce varying outputs even with identical inputs. While we mitigate this by setting the temperature parameter to 0, ensuring deterministic behavior in most components, the Action Model phase presents a unique challenge. During this phase, screenshots of the game are captured, which include Minecraft’s randomly changing, animated background. This unavoidable game feature introduces an element of unpredictability into the Action Model’s processing, as the visual input varies across runs. Further testing is needed to confirm consistency between runs since we were unable to run it multiple times due to resource and time constraints. Furthermore, the limited number of evaluators in the manual analysis phases could introduce bias.
6.3. External Validity
The primary threat to external validity is the limited scope of our evaluation, which focuses solely on crash bugs in Minecraft: Java Edition. This raises concerns about the generalizability of our findings to other types of bugs (e.g., functional bugs, performance issues) and other games. Minecraft, being a sandbox game with a vast and complex environment, may not be representative of all game genres. Furthermore, the specific mechanics and structure of Minecraft: Java Edition, such as the inclusion of an extensive scripting system, might not be comparable to other games. Therefore, while BugCraft demonstrates promising results within the Minecraft context, caution should be exercised when extrapolating these findings to other games or bug types. Further research involving diverse games and bug categories is necessary to establish the broader applicability of our framework. Additionally, our reliance on publicly available bug reports from Mojira might introduce a bias, as these reports may not fully represent the issues encountered by the entire player base.
7. Conclusion
The paper introduces the BugCraft-Bench, a novel benchmark for Minecraft bug reproduction, and BugCraft, an LLM-based framework for end-to-end crash bug reproduction in Minecraft. BugCraft-Bench addresses the dataset gap in game bug reproduction research, offering a curated set of 86 Minecraft crash bug reports for standardized evaluation. BugCraft framework achieves a 30.23% end-to-end success rate in reproducing crash bugs, with its Step Synthesizer demonstrating 66.28% accuracy in plan generation, highlighting LLMs’ potential for automating game bug reproduction. However, limitations in the Action Model, such as the agent getting stuck in loops and poor decision making by the agent were identified, along with Step Synthesizer challenges in command accuracy and detail completeness, especially for newer Minecraft versions.
Moving forward, several key areas warrant further investigation. Future work should focus on enhancing LLM game environment understanding, potentially through collaboration with game developers to create specialized APIs for enhanced in-game movement and interaction, similar to the high-level API Voyager(wang2023voyager, ) uses, called Mineflayer(mineflayer, ). This could also improve bug reproduction across numerous game versions, addressing the limitations of our current vision-based approach. Furthermore, developing comprehensive datasets and robust ground truth verification methods, particularly for non-crash bugs, are crucial. Future benchmarks should incorporate automated or semi-automated ground truth methods to evaluate a wider range of bug types. Assessing BugCraft’s generalizability beyond Minecraft is also essential, requiring investigation into its adaptability to diverse game engines, genres, and complexities through comparative studies. Our roadmap includes improving BugCraft based on identified weaknesses, including addressing challenges with the Action Model and Step Synthesizer. We also plan to extend the agent to handle multiplayer bugs and bugs involving file loading (Worlds, datapacks, world setting JSONs). Ultimately, the advancements presented in this paper provide a strong foundation for future advancements in automated bug reproduction, paving the way for more automatic game development processes.